OpenStudy (anonymous):

PERCENTS: in belvidere, chrylser produced 60 more vehicles this month than last month. if they produced 500 vehicles last month, what was there percent of increase?

OpenStudy (oaktree):

60/500 = 6/50 = 12/100 = 12% increase.